OpenStudy (anonymous):

open parentheses 7 plus square root of 2 close parentheses open parentheses 8 minus 3 square root of 2 close parentheses 50 50 minus 13 square root of 2 50 plus 28 square root of 2 62 minus13 square root of 2

OpenStudy (anonymous):

well,is that it? \[(7+\sqrt{2})(8 - 3\sqrt{2})\]

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

well,then first...do u know how to do this? (a+b)(c-d) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

foil?

OpenStudy (anonymous):

i'll help u: (a+b)(c-d) = a(c-d) + b(c-d) --> (a+b)(c-d) =ac - ad + bc - bd

OpenStudy (anonymous):

and now we have \((7+\sqrt{2})(8 - 3\sqrt{2})\)

OpenStudy (anonymous):

i dont understand. do you multiply whats in the square roots or add them?

OpenStudy (driftracer305):

distribute each term

OpenStudy (driftracer305):

buzz??

OpenStudy (anonymous):